# Decoherence of Histories: Chaotic Versus Integrable Systems

### v1

@@ Abstract
The emergence of decoherence is investigated in Figs. S9 and S10 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. S9 (identical to the dotted line in Fig. S8) and an equilibrium time scale $T_{\text{eq}}$ in Fig. S10 (identical to the dash-dotted line in Fig. S8). Finally, each data point in Figs. S9 and S10 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{28-N}$ different realizations of $|\psi_{0}\rangle$.
@@
For the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D\propto 2^{N}$) with $\alpha\approx 0.3$ at the nonequilibrium time scale and with $\alpha\approx 0.5$ at the equilibrium time scale (with the same $\alpha$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$). This again indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the non-interacting integrable case (ii) we find smaller exponents $\alpha\approx 0.1$ in all cases, yet it is remarkable compared to the free case of the Heisenberg chain studied in the main text that there seemingly is a scaling law. This illustrates how subtle the emergence of decoherence depends on the model under study and a possible reason for this descrepancy could also be the structure of the observable itself instead of the considered model.
The behaviour of decoherence as a function of $T$, is further studied in Fig. S11. For chaotic case (i) $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ consistently decrease and the exponent grows to $\alpha\approx 0.5$ for increasing $T$. In contrast, for integrable case (ii), decoherence becomes (approximately) weaker for larger $T$, but it is again not as clear as for the example of the main text.
To summarize, in the Ising model we find again a strong difference between the chaotic and integrable case in terms of the finite size scaling of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ and their behavior as a function of $T$. This is qualitatively in unison with what we found in the main text and for the chaotic case it is even quantitatively in unison. However, we also found that the free case in the Ising model displays a much stronger form of decoherence compared to the free case in the Heisenberg model, showing that much care is required when one wants to generalize conclusions obtained for one specific integrable model.
@@
The study the chaocity (integrability) of the considered models, we analyze the distribution of the nearest-level spacing of the unfolded spectrum. After unfolding, the averaged level density becomes constant (usually set to $1$, as is done here). The ordered eigenvalues of the unfolded spectrum are denoted by $\tilde{E}_{i}$. As an indicator of quantum chaos, we study the distribution of $s_{i}=\tilde{E}_{i+1}-\tilde{E}_{i}$, denoted by $P(s)$. $P(s)$ differentiates between chaotic and integrable systems: i) For chaotic systems, $P(s)$ follows a Wigner-Dyson distribution, where for systems with time-reversal symmetry,
@@
In both models, due to the existence of additional global symmetries (e.g., translational invariance, reflection, et al.) alongside with the total energy, our analysis is confined to a specific subspace. We compute the $P(s)$ by considering $1/2$ of the total eigenvalues located in the middle of the spectrum, and results are shown in Fig. S12. Good agreement with the Wigner-Dyson distribution is observed in the XXZ model ($\Delta_{1}=1.5,\ \Delta_{2}=0.5$) and in the Ising model ($h_{z}=0.5$), indicating the systems are chaotic with respect to the corresponding parameter. In contrast, in XXZ model ($\Delta_{1}=1.5,\ \Delta_{2}=0.0$), a Possion distribution is found, which suggests that the system is integrable. The results are in line with our findings in the main text. The results for trivial cases, where the model is equivalent to free fermions, e.g., XXZ model ($\Delta_{1}=\Delta_{2}=0$) and Ising model $(h_{z}=0.0)$, are not shown here.

### v2

@@ Abstract
The emergence of decoherence is investigated in Figs. S8 and S9 for Haar random initial states $|\psi_{0}\rangle$. We plot in double logarithmic scale $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus the Hilbert space dimension $D$ for histories of lengths $L\in\{3,4,5\}$ for two different $T$: a nonequilibrium time scale $T_{\text{neq}}$ in Fig. S8 (identical to the dotted line in Fig. S7) and an equilibrium time scale $T_{\text{eq}}$ in Fig. S9 (identical to the dash-dotted line in Fig. S7). Finally, each data point in Figs. S8 and S9 is obtained by averaging $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ over $2^{28-N}$ different realizations of $|\psi_{0}\rangle$.
@@
Figure S11: Average $\overline{\epsilon}$ and maximum $\epsilon_{\text{max}}$ amount of coherence versus relative time step $T/T_{\text{neq}}$ for history length $L=5$ for the (i) chaotic (dark blue) and the (ii) non-interacting integrable (light blue) cases in the Ising model in double logarithmic scale. The system size is $N=22$
For the chaotic case (i) we find a scaling law of the form $D^{-\alpha}$ (note that $D\propto 2^{N}$) with $\alpha\approx 0.25$ at the nonequilibrium time scale and with $\alpha\approx 0.5$ at the equilibrium time scale (with the same $\alpha$ for both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$). This again indicates a robust exponential suppression (with respective to system size $N$) of coherences in chaotic systems. For the non-interacting integrable case (ii) we find smaller exponents $\alpha\approx 0.1$ at the nonequilibrium time scale, and $\alpha\approx 0$ at the equilibrium time scale.
The behaviour of decoherence as a function of $T$ is further studied in Fig. S10. For chaotic case (i) $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ consistently decrease and the exponent grows to $\alpha\approx 0.5$ for increasing $T$. In contrast, for integrable case (ii), decoherence becomes (approximately) weaker for larger $T$, but it is again not as clear as for the example of the main text.
Furthermore, similar to Fig. 4 in the main text, we plot $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ versus $T$ for system size $N=22$ and history length $L=5$ in the Ising model in Fig. S11. Similar trends are observed: in chaotic case, both $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ decrease with increasing $T$. In contrast, in the non-interacting integrable case, they increase on average (after a sudden drop at small $T$), accompanied by significant fluctuations.
To summarize, in the Ising model we find again a strong difference between the chaotic and integrable case in terms of the finite size scaling of $\overline{\epsilon}$ and $\epsilon_{\text{max}}$ and their behavior as a function of $T$. This is qualitatively in unison with what we found in the main text. However, we also found that the free case in the Ising model displays a much stronger form of decoherence compared to the free case in the Heisenberg model, showing that much care is required when one wants to generalize conclusions obtained for one specific integrable model.
@@
To study the chaoticity (integrability) of the considered models, we analyze the distribution of the nearest-level spacing of the unfolded spectrum. After unfolding, the averaged level density becomes constant (usually set to $1$, as is done here). The ordered eigenvalues of the unfolded spectrum are denoted by $\tilde{E}_{i}$. As an indicator of quantum chaos, we study the distribution of $s_{i}=\tilde{E}_{i+1}-\tilde{E}_{i}$, denoted by $P(s)$. $P(s)$ differentiates between chaotic and integrable systems: i) For chaotic systems, $P(s)$ follows a Wigner-Dyson distribution, where for systems with time-reversal symmetry,
@@
In both models, due to the existence of additional global symmetries (translational invariance, reflection invariance, etc.) alongside with total energy conservation, our analysis is confined to a specific subspace. We compute the $P(s)$ by considering $1/2$ of the total eigenvalues located in the middle of the spectrum, and results are shown in Fig. S12. Good agreement with the Wigner-Dyson distribution is observed in the XXZ model ($\Delta_{1}=1.5,\ \Delta_{2}=0.5$) and in the Ising model ($h_{z}=0.5$), indicating the systems are chaotic with respect to the corresponding parameter. In contrast, in XXZ model ($\Delta_{1}=1.5,\ \Delta_{2}=0.0$), a Possion distribution is found, which suggests that the system is integrable. The results are in line with our findings in the main text. The results for trivial cases, where the model is equivalent to free fermions, e.g., XXZ model ($\Delta_{1}=\Delta_{2}=0$) and Ising model $(h_{z}=0.0)$, are not shown here.